OpenStudy (anonymous):

For the given functions f and g, find the requested function. Please show all of your work. f(x) = x+7; g(x) = 6x^2 - x. Find f - g.

OpenStudy (anonymous):

\[f(x)-g(x)=(x+7)-(6x^2-x)\] then algebra

OpenStudy (anonymous):

Can you please explain how you got this? thank you!

OpenStudy (anonymous):

i really just stuck a minus sign between the two functions

OpenStudy (anonymous):

it says to find f and g and Im so confused.lol

OpenStudy (anonymous):

but be careful with the parentheses

OpenStudy (anonymous):

the way i read the problem it GIVES you f and g and asks you to find \[f-g\]

OpenStudy (anonymous):

you wrote f(x) = x+7; g(x) = 6x^2 - x. Find f - g.

OpenStudy (anonymous):

Oh gotcha. Im so confused on this stuff.

OpenStudy (anonymous):

don't forget the distributive law when you remove the parentheses

OpenStudy (anonymous):

you should get \[x+7-(6x^2-x)=x+7-6x^2-x=-6x^2+7\]

OpenStudy (anonymous):

ok so is the first post you put considered the answer, or am i having to simplify further?

OpenStudy (anonymous):

i would write what i wrote at the end. you should do the algebra too so "final answer" would be \[(f-g)(x)=-6x^2+7\]

OpenStudy (anonymous):

i think you are confused on the next one as well yes?

OpenStudy (anonymous):

yes I am! lol- i really appreciate your help